OpenStudy (anonymous):

What is the equation for a line that passes through the points (-3, -9) and (2, 1)

jigglypuff314 (jigglypuff314):

\[\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange} {\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange} {\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\ \Huge\sf\color{blue}{~~~~Welcome ~To~ OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color {blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue} {\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue} {\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\] \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given points (x1, y1) and (x2, y2) so when given the points (-3, -9) and (2, 1) plug in x1=-3 , y1=-9 , x2=2 , y2=1

jigglypuff314 (jigglypuff314):

then once you get your slope, you can plug it in as m for either: slope-intercept form: y = mx + b (where m=slope and b=y-intercept) point-slope form: y - y1 = m(x - x1) (where m=slope and given the point (x1, y1))